\begin{thebibliography}{10}

\bibitem{mtsummit-2021-technologies}
John Ortega, Atul~Kr. Ojha, Katharina Kann, and Chao-Hong Liu, Eds.,
\newblock {\em Proceedings of the 4th Workshop on Technologies for MT of Low Resource Languages (LoResMT2021)}, Virtual, Aug. 2021. Association for Machine Translation in the Americas.

\bibitem{amta-2022-biennial-association}
John~E. Ortega, Marine Carpuat, William Chen, Katharina Kann, Constantine Lignos, Maja Popovic, and Shabnam Tafreshi, Eds.,
\newblock {\em Proceedings of the 15th biennial conference of the Association for Machine Translation in the Americas (Workshop 2: Corpus Generation and Corpus Augmentation for Machine Translation)}. Association for Machine Translation in the Americas, Sept. 2022.

\bibitem{americasnlp-2021-natural}
Manuel Mager, Arturo Oncevay, Annette Rios, Ivan Vladimir~Meza Ruiz, Alexis Palmer, Graham Neubig, and Katharina Kann, Eds.,
\newblock {\em Proceedings of the First Workshop on Natural Language Processing for Indigenous Languages of the Americas}, Online, June 2021. Association for Computational Linguistics.

\bibitem{americasnlp-2023-natural}
Manuel Mager, Abteen Ebrahimi, Arturo Oncevay, Enora Rice, Shruti Rijhwani, Alexis Palmer, and Katharina Kann, Eds.,
\newblock {\em Proceedings of the Workshop on Natural Language Processing for Indigenous Languages of the Americas (AmericasNLP)}, Toronto, Canada, July 2023. Association for Computational Linguistics.

\bibitem{orife2020masakhane}
Iroro Orife, Julia Kreutzer, Blessing Sibanda, Daniel Whitenack, Kathleen Siminyu, Laura Martinus, Jamiil~Toure Ali, Jade Abbott, Vukosi Marivate, Salomon Kabongo, et~al.,
\newblock ``Masakhane--machine translation for africa,''
\newblock {\em arXiv preprint arXiv:2003.11529}, 2020.

\bibitem{nekoto-etal-2020-participatory}
Wilhelmina Nekoto, Vukosi Marivate, Tshinondiwa Matsila, Timi Fasubaa, Taiwo Fagbohungbe, Solomon~Oluwole Akinola, Shamsuddeen Muhammad, Salomon Kabongo~Kabenamualu, Salomey Osei, Freshia Sackey, Rubungo~Andre Niyongabo, Ricky Macharm, Perez Ogayo, Orevaoghene Ahia, Musie~Meressa Berhe, Mofetoluwa Adeyemi, Masabata Mokgesi-Selinga, Lawrence Okegbemi, Laura Martinus, Kolawole Tajudeen, Kevin Degila, Kelechi Ogueji, Kathleen Siminyu, Julia Kreutzer, Jason Webster, Jamiil~Toure Ali, Jade Abbott, Iroro Orife, Ignatius Ezeani, Idris~Abdulkadir Dangana, Herman Kamper, Hady Elsahar, Goodness Duru, Ghollah Kioko, Murhabazi Espoir, Elan van Biljon, Daniel Whitenack, Christopher Onyefuluchi, Chris~Chinenye Emezue, Bonaventure F.~P. Dossou, Blessing Sibanda, Blessing Bassey, Ayodele Olabiyi, Arshath Ramkilowan, Alp {\"O}ktem, Adewale Akinfaderin, and Abdallah Bashir,
\newblock ``Participatory research for low-resourced machine translation: A case study in {A}frican languages,''
\newblock in {\em Findings of the Association for Computational Linguistics: EMNLP 2020}, Online, Nov. 2020, pp. 2144--2160, Association for Computational Linguistics.

\bibitem{ebrahimi-etal-2023-findings}
Abteen Ebrahimi, Manuel Mager, Shruti Rijhwani, Enora Rice, Arturo Oncevay, Claudia Baltazar, Mar{\'\i}a Cort{\'e}s, Cynthia Monta{\~n}o, John~E. Ortega, Rolando Coto-solano, Hilaria Cruz, Alexis Palmer, and Katharina Kann,
\newblock ``Findings of the {A}mericas{NLP} 2023 shared task on machine translation into indigenous languages,''
\newblock in {\em Proceedings of the Workshop on Natural Language Processing for Indigenous Languages of the Americas (AmericasNLP)}, Toronto, Canada, July 2023, pp. 206--219, Association for Computational Linguistics.

\bibitem{mager-etal-2021-findings}
Manuel Mager, Arturo Oncevay, Abteen Ebrahimi, John Ortega, Annette Rios, Angela Fan, Ximena Gutierrez-Vasques, Luis Chiruzzo, Gustavo Gim{\'e}nez-Lugo, Ricardo Ramos, Ivan~Vladimir Meza~Ruiz, Rolando Coto-Solano, Alexis Palmer, Elisabeth Mager-Hois, Vishrav Chaudhary, Graham Neubig, Ngoc~Thang Vu, and Katharina Kann,
\newblock ``Findings of the {A}mericas{NLP} 2021 shared task on open machine translation for indigenous languages of the {A}mericas,''
\newblock in {\em Proceedings of the First Workshop on Natural Language Processing for Indigenous Languages of the Americas}, Online, June 2021, pp. 202--217, Association for Computational Linguistics.

\bibitem{chen2023reducing}
William Chen, Xuankai Chang, Yifan Peng, Zhaoheng Ni, Soumi Maiti, and Shinji Watanabe,
\newblock ``Reducing barriers to self-supervised learning: Hubert pre-training with academic compute,''
\newblock in {\em Proc. Interspeech 2023}, 2023.

\bibitem{babu2021xls}
Arun Babu, Changhan Wang, Andros Tjandra, Kushal Lakhotia, Qiantong Xu, Naman Goyal, Kritika Singh, Patrick von Platen, Yatharth Saraf, Juan Pino, et~al.,
\newblock ``{XLS-R}: Self-supervised cross-lingual speech representation learning at scale,''
\newblock {\em arXiv preprint arXiv:2111.09296}, 2021.

\bibitem{radford2022robust}
Alec Radford, Jong~Wook Kim, Tao Xu, Greg Brockman, Christine McLeavey, and Ilya Sutskever,
\newblock ``Robust speech recognition via large-scale weak supervision,''
\newblock {\em arXiv preprint arXiv:2212.04356}, 2022.

\bibitem{pratap20_interspeech_mls}
Vineel Pratap, Qiantong Xu, Anuroop Sriram, Gabriel Synnaeve, and Ronan Collobert,
\newblock ``{MLS: A Large-Scale Multilingual Dataset for Speech Research},''
\newblock in {\em Proc. Interspeech 2020}, 2020, pp. 2757--2761.

\bibitem{zhang2023google}
Yu~Zhang, Wei Han, James Qin, Yongqiang Wang, Ankur Bapna, Zhehuai Chen, Nanxin Chen, Bo~Li, Vera Axelrod, Gary Wang, et~al.,
\newblock ``Google {USM}: Scaling automatic speech recognition beyond 100 languages,''
\newblock {\em arXiv preprint arXiv:2303.01037}, 2023.

\bibitem{conneau2022fleurs}
Alexis Conneau, Min Ma, Simran Khanuja, Yu~Zhang, Vera Axelrod, Siddharth Dalmia, Jason Riesa, Clara Rivera, and Ankur Bapna,
\newblock ``{FLEURS}: Few-shot learning evaluation of universal representations of speech,''
\newblock in {\em 2022 IEEE Spoken Language Technology Workshop (SLT)}, 2023, pp. 798--805.

\bibitem{shi2023exploration}
Jiatong Shi, Yun Tang, Hirofumi Inaguma, Hongyu GOng, Juan Pino, and Shinji Watanabe,
\newblock ``Exploration on {HuBERT} with multiple resolutions,''
\newblock {\em arXiv preprint arXiv:2306.01084}, 2023.

\bibitem{gales2014speech}
Mark~JF Gales, Kate~M Knill, Anton Ragni, and Shakti~P Rath,
\newblock ``Speech recognition and keyword spotting for low-resource languages: Babel project research at cued,''
\newblock in {\em Fourth International workshop on spoken language technologies for under-resourced languages (SLTU-2014)}. International Speech Communication Association (ISCA), 2014, pp. 16--23.

\bibitem{agrawal-etal-2023-findings}
Milind Agarwal, Sweta Agrawal, Antonios Anastasopoulos, Luisa Bentivogli, Ond{\v{r}}ej Bojar, Claudia Borg, Marine Carpuat, Roldano Cattoni, Mauro Cettolo, Mingda Chen, William Chen, Khalid Choukri, Alexandra Chronopoulou, Anna Currey, Thierry Declerck, Qianqian Dong, Kevin Duh, Yannick Est{\`e}ve, Marcello Federico, Souhir Gahbiche, Barry Haddow, Benjamin Hsu, Phu Mon~Htut, Hirofumi Inaguma, D{\'a}vid Javorsk{\'y}, John Judge, Yasumasa Kano, Tom Ko, Rishu Kumar, Pengwei Li, Xutai Ma, Prashant Mathur, Evgeny Matusov, Paul McNamee, John P.~McCrae, Kenton Murray, Maria Nadejde, Satoshi Nakamura, Matteo Negri, Ha~Nguyen, Jan Niehues, Xing Niu, Atul Kr.~Ojha, John E.~Ortega, Proyag Pal, Juan Pino, Lonneke van~der Plas, Peter Pol{\'a}k, Elijah Rippeth, Elizabeth Salesky, Jiatong Shi, Matthias Sperber, Sebastian St{\"u}ker, Katsuhito Sudoh, Yun Tang, Brian Thompson, Kevin Tran, Marco Turchi, Alex Waibel, Mingxuan Wang, Shinji Watanabe, and Rodolfo Zevallos,
\newblock ``{FINDINGS} {OF} {THE} {IWSLT} 2023 {EVALUATION} {CAMPAIGN},''
\newblock in {\em Proceedings of the 20th International Conference on Spoken Language Translation (IWSLT 2023)}, Toronto, Canada (in-person and online), July 2023, pp. 1--61, Association for Computational Linguistics.

\bibitem{rios-gonzales-castro-mamani-2014-morphological}
Annette Rios~Gonzales and Richard~Alexander Castro~Mamani,
\newblock ``Morphological disambiguation and text normalization for {S}outhern {Q}uechua varieties,''
\newblock in {\em Proceedings of the First Workshop on Applying {NLP} Tools to Similar Languages, Varieties and Dialects}, Dublin, Ireland, Aug. 2014, pp. 39--47, Association for Computational Linguistics and Dublin City University.

\bibitem{rios2010applying}
Annette Rios,
\newblock ``Applying finite-state techniques to a native american language: Quechua,''
\newblock .

\bibitem{rios-gonzales-gohring-2013-machine}
Annette Rios~Gonzales and Anne G{\"o}hring,
\newblock ``Machine learning disambiguation of {Q}uechua verb morphology,''
\newblock in {\em Proceedings of the Second Workshop on Hybrid Approaches to Translation}, Sofia, Bulgaria, Aug. 2013, pp. 13--18, Association for Computational Linguistics.

\bibitem{rios2015basic}
Annette Rios,
\newblock {\em A basic language technology toolkit for quechua},
\newblock Ph.D. thesis, University of Zurich, 2015.

\bibitem{rios2008quechua}
Annette Rios, Anne G{\"o}hring, and Martin Volk,
\newblock ``A quechua-spanish parallel treebank,''
\newblock {\em Lot occasional series}, vol. 12, pp. 53--64, 2008.

\bibitem{rios2011spell}
``Spell checking an agglutinative language: Quechua,''
\newblock 2011.

\bibitem{ortega-pillaipakkamnatt-2018-using}
John Ortega and Krishnan Pillaipakkamnatt,
\newblock ``Using morphemes from agglutinative languages like {Q}uechua and {F}innish to aid in low-resource translation,''
\newblock in {\em Proceedings of the {AMTA} 2018 Workshop on Technologies for {MT} of Low Resource Languages ({L}o{R}es{MT} 2018)}, Boston, MA, Mar. 2018, pp. 1--11, Association for Machine Translation in the Americas.

\bibitem{ortega2020neural}
John~E Ortega, Richard Castro~Mamani, and Kyunghyun Cho,
\newblock ``Neural machine translation with a polysynthetic low resource language,''
\newblock {\em Machine Translation}, vol. 34, no. 4, pp. 325--346, 2020.

\bibitem{chen-fazio-2021-morphologically}
William Chen and Brett Fazio,
\newblock ``Morphologically-guided segmentation for translation of agglutinative low-resource languages,''
\newblock in {\em Proceedings of the 4th Workshop on Technologies for MT of Low Resource Languages (LoResMT2021)}, Virtual, Aug. 2021, pp. 20--31, Association for Machine Translation in the Americas.

\bibitem{zevallos-etal-2022-introducing}
Rodolfo Zevallos, John Ortega, William Chen, Richard Castro, N{\'u}ria Bel, Cesar Toshio, Renzo Venturas, Hilario Aradiel, and Nelsi Melgarejo,
\newblock ``Introducing {Q}u{BERT}: A large monolingual corpus and {BERT} model for {S}outhern {Q}uechua,''
\newblock in {\em Proceedings of the Third Workshop on Deep Learning for Low-Resource Natural Language Processing}, Hybrid, July 2022, pp. 1--13, Association for Computational Linguistics.

\bibitem{cardenas2018siminchik}
Ronald Cardenas, Rodolfo Zevallos, Reynaldo Baquerizo, and Luis Camacho,
\newblock ``Siminchik: A speech corpus for preservation of southern quechua,''
\newblock {\em ISI-NLP 2}, p.~21, 2018.

\bibitem{zevallos-etal-2022-huqariq}
Rodolfo Zevallos, Luis Camacho, and Nelsi Melgarejo,
\newblock ``Huqariq: A multilingual speech corpus of native languages of {P}eru for{S}peech recognition,''
\newblock in {\em Proceedings of the Thirteenth Language Resources and Evaluation Conference}, Marseille, France, June 2022, pp. 5029--5034, European Language Resources Association.

\bibitem{vaswani2017attention}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan~N Gomez, {\L}ukasz Kaiser, and Illia Polosukhin,
\newblock ``Attention is all you need,''
\newblock {\em Proc. NeurIPS}, vol. 30, 2017.

\bibitem{e-ortega-etal-2023-quespa}
John E.~Ortega, Rodolfo Zevallos, and William Chen,
\newblock ``{QUESPA} submission for the {IWSLT} 2023 dialect and low-resource speech translation tasks,''
\newblock in {\em Proceedings of the 20th International Conference on Spoken Language Translation (IWSLT 2023)}, Toronto, Canada (in-person and online), July 2023, pp. 261--268, Association for Computational Linguistics.

\bibitem{conneau2020unsupervised}
Alexis Conneau, Alexei Baevski, Ronan Collobert, Abdelrahman Mohamed, and Michael Auli,
\newblock ``Unsupervised cross-lingual representation learning for speech recognition,''
\newblock {\em arXiv preprint arXiv:2006.13979}, 2020.

\bibitem{giannakopoulos2015pyaudioanalysis}
Theodoros Giannakopoulos,
\newblock ``pyaudioanalysis: An open-source python library for audio signal analysis,''
\newblock {\em PloS one}, vol. 10, no. 12, pp. e0144610, 2015.

\bibitem{shi2023ml}
Jiatong Shi, Dan Berrebbi, William Chen, Ho-Lam Chung, En-Pei Hu, Wei~Ping Huang, Xuankai Chang, Shang-Wen Li, Abdelrahman Mohamed, Hung-yi Lee, et~al.,
\newblock ``{ML-SUPERB}: {M}ulti{L}ingual {S}peech {U}niversal {PER}formance {B}enchmark,''
\newblock in {\em Proc. Interspeech 2023}, 2023.

\bibitem{schneider19_wav2vec}
Steffen Schneider, Alexei Baevski, Ronan Collobert, and Michael Auli,
\newblock ``{wav2vec: Unsupervised Pre-Training for Speech Recognition},''
\newblock in {\em Proc. Interspeech 2019}, 2019, pp. 3465--3469.

\bibitem{baevskiw2v}
Alexei Baevski, Yuhao Zhou, Abdelrahman Mohamed, and Michael Auli,
\newblock ``{wav2vec} 2.0: A framework for self-supervised learning of speech representations,''
\newblock in {\em Advances in Neural Information Processing Systems}. 2020, vol.~33, pp. 12449--12460, Curran Associates, Inc.

\bibitem{lee-etal-2022-textless}
Ann Lee, Hongyu Gong, Paul-Ambroise Duquenne, Holger Schwenk, Peng-Jen Chen, Changhan Wang, Sravya Popuri, Yossi Adi, Juan Pino, Jiatao Gu, and Wei-Ning Hsu,
\newblock ``Textless speech-to-speech translation on real data,''
\newblock in {\em Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies}, Seattle, United States, July 2022, pp. 860--872, Association for Computational Linguistics.

\bibitem{hsuHubert}
Wei-Ning Hsu, Benjamin Bolte, Yao-Hung~Hubert Tsai, Kushal Lakhotia, Ruslan Salakhutdinov, and Abdelrahman Mohamed,
\newblock ``{HuBERT}: Self-supervised speech representation learning by masked prediction of hidden units,''
\newblock {\em IEEE/ACM Transactions on Audio, Speech, and Language Processing}, vol. 29, pp. 3451--3460, 2021.

\bibitem{watanabe2018espnet}
Shinji Watanabe, Takaaki Hori, Shigeki Karita, Tomoki Hayashi, Jiro Nishitoba, Yuya Unno, Nelson {Enrique Yalta Soplin}, Jahn Heymann, Matthew Wiesner, Nanxin Chen, Adithya Renduchintala, and Tsubasa Ochiai,
\newblock ``{ESPnet}: End-to-end speech processing toolkit,''
\newblock in {\em Proceedings of Interspeech}, 2018, pp. 2207--2211.

\bibitem{park2019specaugment}
Daniel~S Park, William Chan, Yu~Zhang, Chung-Cheng Chiu, Barret Zoph, Ekin~D Cubuk, and Quoc~V Le,
\newblock ``Specaugment: A simple data augmentation method for automatic speech recognition,''
\newblock {\em Proc. Interspeech}, pp. 2613--2617, 2019.

\bibitem{graves2006connectionist}
Alex Graves, Santiago Fern{\'a}ndez, Faustino Gomez, and J{\"u}rgen Schmidhuber,
\newblock ``Connectionist temporal classification: labelling unsegmented sequence data with recurrent neural networks,''
\newblock in {\em Proc. International Conference on Machine Learning}, 2006, pp. 369--376.

\bibitem{kingma2014adam}
Diederik~P Kingma and Jimmy Ba,
\newblock ``Adam: A method for stochastic optimization,''
\newblock in {\em Proc. ICLR}, 2015.

\end{thebibliography}

\documentclass{article}
\usepackage{spconf,amsmath,graphicx}
\usepackage{multicol}
\usepackage{multirow}
\usepackage{arydshln}
\usepackage{xcolor}
\usepackage{lipsum}
\usepackage{microtype}
\usepackage{graphicx}
\usepackage{subfigure}
\usepackage[hidelinks]{hyperref}       
\usepackage{booktabs} 
\usepackage{amsfonts}       
\usepackage{nicefrac}       
\usepackage{arydshln}
\usepackage{algpseudocode}
\usepackage{algorithm}

\usepackage{tikz}
\usepackage{pgfplots}
\title{Evaluating Self-Supervised Speech Representations for Indigenous American Languages}
\name{Chih-Chen Chen$^1$, William Chen$^2$, Rodolfo Zevallos$^3$, John E. Ortega$^4$}
\address{Taipei Medical University$^1$ \\Carnegie Mellon University$^2$ \\  Universitat Pompeu Fabra$^3$ \\ Northeastern University$^4$}
\begin{document}
%
\maketitle
\begin{abstract}
The application of self-supervision to speech representation learning has garnered significant interest in recent years, due to its scalability to large amounts of unlabeled data. However, much progress, both in terms of pre-training and downstream evaluation, has remained concentrated in monolingual models that only consider English. Few models consider other languages, and even fewer consider indigenous ones. In our submission to the New Language Track of the ASRU 2023 ML-SUPERB Challenge, we present an ASR corpus for Quechua, an indigenous South American Language. We benchmark the efficacy of large SSL models on Quechua, along with 6 other indigenous languages such as Guarani and Bribri, on low-resource ASR. Our results show surprisingly strong performance by state-of-the-art SSL models, showing the potential generalizability of large-scale models to real-world data.
\end{abstract}
\begin{keywords}
Indigenous Languages, Low-resource, ML-SUPERB, Self-Supervised Learning
\end{keywords}
\section{Introduction}
\label{sec:intro}
In recent years, the fields of Natural Language Processing (NLP) and speech processing has witnessed remarkable advancements, with applications ranging from machine translation and sentiment analysis to voice assistants and chatbots. These developments have predominantly focused on widely spoken languages such as English and Mandarin. However, the vast linguistic diversity represented by indigenous languages across the globe remains largely unexplored in the context of language processing.

Indigenous languages are the ancestral tongues of diverse communities with rich cultural heritage and profound connections to the environment in which they are spoken. These languages often exhibit distinct linguistic characteristics, deviating from the structures and conventions of widely studied languages. By expanding the scope of language processing to include indigenous languages, we can foster linguistic inclusivity and empower indigenous communities to participate in the digital era while preserving their linguistic and cultural identities.

One compelling reason to focus on indigenous languages in language processing is the potential for social impact. Many Indigenous communities face challenges related to limited access to information and technology, which further exacerbate social and economic disparities. By developing NLP models and applications tailored to indigenous languages, we can bridge the digital divide and enable these communities to leverage technology for communication, education, and cultural preservation. Such efforts have the potential to enhance language revitalization efforts, foster inter-generational transmission of knowledge, and promote cultural preservation within indigenous communities.

Our core contribution is a Quechua ASR corpus, which serves as our submission to the New Language Track of the ML-SUPERB Challenge. We evaluate the effectiveness of different self-supervised learning (SSL) models for speech on this corpus, along with other indigenous American languages: Bribri, Guarani, Kotiria, Wa'ikhana, and Totonac.

We first introduce general characteristics of American indigenous languages and discuss the challenges in modelling them due to their unique linguistic natures. We then provide a brief overview of each language to this study, highlighting some key linguistic properties. As our other core contribution, we discuss research in American indigenous languages in both the fields of NLP and speech processing, hoping to create a bridge in the literature for communities. 

\section{American Indigenous Languages}

American Indigenous Languages encompass a diverse range of language families and isolates, each with its own linguistic features. The languages of this region exhibit a remarkable variety of phonological, morphological, syntactic, and semantic structures, reflecting the rich linguistic diversity of the continent. A persistent challenge in modelling these languages, similar to many indigenous languages, is the frequency of code-switching. Coupled with the lack of both linguistic and electronic resources for these languages, creating language technologies for indigenous languages remains a significant challenge despite the exponential progress in NLP and speech processing.

Broadly speaking, the indigenous languages of the Americas are morphologically-rich, often exhibiting agglutinative or polysynthetic structures. These languages tend to have extensive systems of affixation, where morphemes are added to roots to convey meaning and grammatical information. For example, in Quechua or Guarani, complex words can be formed through the addition of numerous affixes to a single root. This makes them particularly challenging for NLP and language modelling tasks, due to the higher frequency of rare words.

\subsection{Quechua}
Quechua is a family of closely related languages spoken by around 10 million people across South America. While primarily spoken in the Andean regions, communities can also be found along the plains and valleys connecting the Amazon to the Pacific. Quechua is considered one of the most widely spoken indigenous language families in the Americas. While there are regional variations, Quechua languages share many common linguistic characteristics. These variations are broadly separated into two distinct categories: Quechua I and Quechua II. The former refers to the varieties of Quechua spoken in the central parts of Peru, while the latter is spoken in Southern Peru, Bolivia, and Colombia. Dialects between the two categories may not necessarily be mutually intelligible, which is what distinguishes Quechua as a language family, and makes it particularly challenging for language processing.

\subsection{Bribri}
Bribri, also known as the Bribri-Poró language, is spoken by the Bribri people of Costa Rica. It belongs to the Chibchan language family, which is primarily found in Central America. The Bribri language specifically falls under the Guaymí subgroup of the Chibchan family. Geographically, the Bribri language is primarily spoken in the Talamanca region of Costa Rica, specifically in the southern parts of Limón and northern parts of Puntarenas provinces. It is a tonal language, meaning that pitch variations can distinguish between different words or meanings. The language also exhibits agglutinative tendencies, where words are formed by adding affixes to a base or root.

\subsection{Guarani}
The Guarani language is an indigenous language spoken by the Guarani people in South America. It is a member of the Tupi-Guarani language family, which encompasses several languages across Brazil, Paraguay, Argentina, and Bolivia. Guarani is one of the most widely spoken indigenous languages in the Americas, with 4-6 million speakers. Guarani is mainly distributed in Paraguay, where it has official status alongside Spanish. It is also spoken in parts of northeastern Argentina, southeastern Bolivia, and southern Brazil. Like many other American languages, Guarani is agglutinative. Guarani is a highly regular language with consistent pronunciation rules, which may be notable for speech processing.

\subsection{Kotiria}
Kotiria, also known as Wanano, is an indigenous language spoken by the Kotiria people who reside in the Vaupés region of Colombia and Brazil. Kotiria is part of the larger Eastern Tukanoan language family, which includes several other indigenous languages spoken in the northwest Amazon region. Kotiria language is primarily spoken in the upper and middle basins of the Vaupés River, which runs through the Amazon rainforest. The language is concentrated in remote areas of the Colombian Vaupés Department and the Brazilian state of Amazonas. Like Bribri, it is both agglutinative and tonal. 

\subsection{Wa'ikhana}
Also known as Cubeo, Wa'ikhana is spoken by the Cubeo people in the northwest Amazon region, primarily in Colombia and Brazil. Wa'ikhana belongs to the Tucanoan language family, which encompasses several indigenous languages spoken in the northwest Amazon. Wa'ikhana is also both agglutinative and tonal.

\subsection{Totonac}
The Totonac language is an indigenous language spoken by the Totonac people in Mexico. It belongs to the Totonacan language family, which is primarily found in the states of Veracruz, Puebla, and parts of Hidalgo in eastern Mexico. Totonac is both agglutinative and tonal.

\section{Indigenous Languages in Language Processing}

\subsection{Community Efforts}

In the field of NLP, several initiatives have been started to encourage further research in indigenous languages. While the majority are workshops for general low-resource NLP \cite{mtsummit-2021-technologies, amta-2022-biennial-association}, newer efforts have also targeted indigenous languages \cite{americasnlp-2021-natural, americasnlp-2023-natural, orife2020masakhane, nekoto-etal-2020-participatory}. For American indigenous languages specifically, the AmericasNLP \cite{americasnlp-2021-natural, americasnlp-2023-natural} community has helped driven research by improving the visibility of authors from indigenous communities. AmericasNLP also hosts an annual shared task, similar to those found in machine and speech translation workshops \cite{ebrahimi-etal-2023-findings, mager-etal-2021-findings}, to further integrate state-of-the-art methods with indigenous languages.

In speech processing, research for indigenous languages is more \textit{ad hoc}, with numerous decentralized efforts from a variety of research groups. Contrary to NLP, indigenous languages play a more common role in SOTA models \cite{chen2023reducing, babu2021xls, radford2022robust, pratap20_interspeech_mls, zhang2023google} and benchmarks \cite{conneau2022fleurs, shi2023exploration, gales2014speech}. Annual challenges, primarily for speech translation, also help bring SOTA methods to these languages \cite{agrawal-etal-2023-findings}.

\subsection{Research for Quechua}
Quechua has enjoyed high amounts of attention in the field of NLP relative to other American languages, likely due to its higher population and speakers and thus larger amount of resources. Much of the early work involved analyzing the morphological properties of Quechua, such as through finite state transducers, \cite{rios-gonzales-castro-mamani-2014-morphological, rios2010applying, rios-gonzales-gohring-2013-machine} and developing toolkits to process the language \cite{rios2015basic, rios2008quechua, rios2011spell}. Neural approaches were first adopted in the context of machine translation \cite{ortega-pillaipakkamnatt-2018-using, ortega2020neural, chen-fazio-2021-morphologically} before being adopted to masked language models \cite{zevallos-etal-2022-introducing}.

Fewer studies have been done on Quechua on the speech processing side, likely due to the lack of available data. To our knowledge, Siminchik \cite{cardenas2018siminchik} was the first speech corpus for Quechua. While the authors established baseline results with HMM-based systems, a full release of the data was never realized. Similarly, Huqariq \cite{zevallos-etal-2022-huqariq} is a multilingual collection of four native Peruvian languages, including Quechua, that has yet to be publicly released. Quechua was featured in two speech processing challenges, the AmericasNLP 2022 Competition and IWSLT 2023 \cite{agrawal-etal-2023-findings}. The latter case saw the first evaluation of Quechua with Transformer-based \cite{vaswani2017attention}, with participants greatly leveraging pre-trained SSL models \cite{e-ortega-etal-2023-quespa}. However, the performance of other SSL methods on Quechua remain an open question, as all participants utilized either XLSR 53 \cite{conneau2020unsupervised} or XLS-R 128 \cite{babu2021xls}.

\section{Corpus} \label{sec:data}

Our submission to the ASRU 2023 ML-SUPERB challenge is derived from the Siminchik corpus \cite{cardenas2018siminchik}, which contains recordings of two different Quechua II dialects. The first is Chanca Quechua, which is spoken primarily in Ayacucho and its surrounding areas in Peru. The other dialect is Collao Quechua, which is spoken in Cusco and Puno.  

Siminchik \cite{cardenas2018siminchik} consists of crowd-resourced transcriptions of radio recordings. The initial recordings were collected from radio channels in Ayachuco and Apurimac for Chanca Quechua, while for Collao Quechua they were collected from Puno and Cusco. Advertisements, music, and segments of Spanish speech were filtered out, yielding 97 hours of audio. Audio clips were then segmented to a maximum of 30 seconds, although beginning and end words were likely trimmed due to the imperfect truncation method \cite{cardenas2018siminchik, giannakopoulos2015pyaudioanalysis}. 

The authors of Siminchik conducted several post-processing steps on the annotated transcripts. First, punctuation was removed and casing was normalized. Due to differences in dialects, interjections were also normalized. This was accomplished by making a dictionary that mapped expressions to a specific word form. Furthermore, dialectal differences also require the spelling of the ASR transcripts to be normalized to a common form. 
This was accomplished using a finite state transducer-based normalization toolkit for Quechua \cite{rios-gonzales-castro-mamani-2014-morphological}, which adheres to the spelling of the Chanca dialect. 

For our submission to ML-SUPERB, we sample 90 minutes of speech from the corpus, to create the 10-minute and 1-hour training set, the 10-minute development set, and the 10-minute test. 

\section{Experimental Setup}

\subsection{Data}

We conduct our experiments on 6 indigenous American languages: Quechua, Bribri, Guarani, Kotiria, Wa'ikhana, and Totonac. For Quechua, we use the data we submitted to the ML-SUPERB Challenge described in Section \ref{sec:data}. For Bribri, Guarani, Kotiria, and Wa'ikhana, we use data from the 2022 AmericasNLP competition. The Totonac data is derived from a study on using speech technologies for endangered languages documentation. Due to the lack of a public testing split for AmericasNLP, we split the data ourselves from the provided validation sets. All data is formatted in the style of ML-SUPERB \cite{shi2023ml}, which consists of a 1-hour training set, a 10-minute training set, a 10-minute validation set, and a 10-minute testing set. This low-resource data setting most accurately benchmarks the capabilities of SSL models for indigenous languages in real-world settings, due to the lack of both available labeled and unlabeled data.
\begin{table*}[htb]
    \centering
    \caption{Evaluation of SSL models on each indigenous language on the 10-minute set, measured in character error rate (CER $\downarrow$).}
    \resizebox {\linewidth} {!} {
\begin{tabular}{lc|cccccc|c}
\toprule
Model & Hours & Quechua & Bribri & Guarani & Kotiria & Wa'ikhana & Totonac & Average \\
\midrule
XLSR 53 & 56k  & 47.8 & 54.6 & 37.6 & 64.2 & 83.3 & 29.6 & 52.9\\
XLS-R 128  & 436k & \textbf{42.5} & \textbf{49.5} & \textbf{27.5} & \textbf{51.2} & \textbf{62.2} & \textbf{27.7} & 43.4\\
mHuBERT & 13.5k & 47.7& 54.3 & 35.2 & 64.8 & 84.8 & 30.1 & 52.8 \\ 
\bottomrule
\end{tabular}
}
    \label{tab:results_10m}
    \vspace{-0.5cm}
\end{table*}

\begin{table*}[htb]
    \centering
    \caption{Evaluation of SSL models on each indigenous language on the 1 hour set, measured in character error rate (CER $\downarrow$).}
    \resizebox {\linewidth} {!} {
\begin{tabular}{lc|cccccc|c}
\toprule
Model & Hours & Quechua & Bribri & Guarani & Kotiria & Wa'ikhana & Totonac & Average \\
\midrule
XLSR 53 & 56k  & 37.5 & 49.5 & 31.5 & 49.9 & 62.4 & 26.0 & 42.8\\
XLS-R 128 & 436k & \textbf{34.0} & \textbf{44.1} & \textbf{24.0} &\textbf{ 43.4 }& \textbf{55.1 }& \textbf{20.6 }& \textbf{36.8} \\
mHuBERT& 13.5k & 37.1 & 49.2 & 32.0 & 50.6 & 62.3 & 26.1 & 42.8\\ 
\bottomrule
\end{tabular}
}
    \label{tab:results_1h}
    \vspace{-0.5cm}
\end{table*}
\subsection{Self-Supervised Models}
We evaluate three SSL models on each language, along with log-Mel filterbank features (FBANK). The models are described as follows:

\subsubsection{XLSR 53}
XLSR 53 \cite{conneau2020unsupervised} is trained on 56k hours of multilingual data for 53 languages, which are pre-dominantly European. It uses the 317M parameter wav2vec architecture \cite{schneider19_wav2vec}, which consists of a convolutional feature extractor and Transformer encoder \cite{vaswani2017attention} trained with contrastive loss.

\subsubsection{XLS-R 128}
XLS-R 128 \cite{babu2021xls} is the large-scale extension of XLSR 53, trained on 436k hours of multilingual data across 128 languages. It instead uses the wav2vec 2.0 \cite{baevskiw2v} architecture, which also includes a convolutional feature extractor and Transformer encoder, but is trained with both contrastive and codebook prediction losss.

\subsubsection{mHuBERT}
mHuBERT \cite{lee-etal-2022-textless} builds off of the HuBERT \cite{hsuHubert} architecture, which uses an iterative approach to SSL. HuBERT models are trained to predict discrete representations of masked speech. After each iteration of pre-training, hidden representations are extracted from the model and clustered using k-means, creating the discrete targets for the next round of pre-training. mHuBERT was trained multilingually on 3 languages: Spanish, French, and Italian, each 4.5k hours of data. It uses the 95M parameter HuBERT Base architecture, which modifies the wav2vec 2.0 design for pure codebook prediction.

\subsection{Training Settings}
We conduct all experiments using the ESPnet \cite{watanabe2018espnet} toolkit with the official settings of the ML-SUPERB competition. The SSL model is used as a frozen feature extractor, such that the hidden representation of each layer is obtained. The layer-wise outputs of combined via a weighted sum, where the weight is learned during training. These outputs are then down projected to a hidden size of 80 and then augmented with SpecAug \cite{park2019specaugment}, before being used as the model inputs of a Transformer encoder \cite{vaswani2017attention}. The Transformer consists of 2 layers, each with a hidden size of 256, 8 attention heads, and a feed-forward size of 1024. Models are trained with CTC loss \cite{graves2006connectionist} and the Adam optimizer \cite{kingma2014adam}, with a constant learning rate of 0.0001. Models are trained for a maximum of 15,000 steps and the 5 best checkpoints are averaged for inference, which is performed with CTC greedy decoding.

\section{Results}

Our experimental results are presented in Tables \ref{tab:results_10m} and \ref{tab:results_1h} for the 10-minute and 1-hour settings respectively. Models are evaluated in character error rate (CER).

Similar to the complete ML-SUPERB benchmark, XLS-R 128 \cite{babu2021xls} obtains the highest overall scores in both data settings. The results presented here are even more distinct: XLS-R 128 outperforms all other models every single task. This suggests the powerful generalizability of large-scale multilingual SSL: all evaluated languages (aside from Guarani) were unseen during pre-training. 

The distance between the other two models, XLSR 53 and mHuBERT, is much smaller, with only a difference of 0.1 average CER on the 10-minute track and no significant difference on the 1-hour track. A strong future research question would be to isolate the cause for the lack of difference, as one would expect the model trained on more languages to generalize better.

Overall, we find the results of our evaluation surprisingly strong. The average CER of XLS-R 128 on the 1 hour set is 36.8, only 6.2 CER higher than its average monolingual score on ML-SUPERB. While further improvements are necessary, this shows that it is possible to shrink the gap between high and low-resource languages with powerful cross-lingual transfer learning, even in cases without large-scale unlabeled data.

\section{Conclusion}
While the recent progress of deep learning in NLP and speech processing has significantly accelerated the development of language technologies, the progress has been unequally distributed. In our submission to the ASRU 2023 ML-SUPERB Challenge, we present an ASR corpus for Quechua, an indigenous South American Language. We are the first benchmark the effectiveness of large-scale speech SSL models on indigenous American languages such as Quechua, which are known to be among the most difficult for NLP. We find surprisingly positive results, showing the impressive generalization ability of large-scale multilingual SSL models on new languages.

\clearpage
\bibliographystyle{IEEEbib}
\bibliography{strings,refs}

\end{document}